OpenStudy (australopithecus):

A truck on a straight road starts from rest, accelerating at 2.00 m/s2 until it reaches a speed of 20.0 m/s. Then the truck travels for 20.0 s at constant speed until the brakes are applied, stopping the truck in a uniform manner in an additional 5.00 s. (a) How long is the truck in motion? (b) What is the average velocity of the truck for the motion described? a) I use the equation V_xf = V_xi + a_x(t) I know V_xf = 20.0m/s V_xi = 0m/s ax = 2.0 thus I solve for t and get 10seconds and then take 5 seconds to break after 20 seconds of driving so for a) 10 + 20 + 5 = 35 seconds

OpenStudy (australopithecus):

for part b) I need to use the formula Delta x/Delta t = average velocity Need to find xf, knowing xi = 0 Thus I use the formula x_f = x_i + V_xi(t) + (1/2)a_x(t)^(2) x_f = 0 + 0(10) + (1/2)(2.0)(10)^(2) x_f = 100m so for the first 10 seconds the truck traveled 100ms At a speed of 20m/s 20m/s = xm/20s 20*20 = x x = 400 thus we have 100+400 = 500m then it slows down from 500m to x_f thus I use the equation x_f = x_i + (1/2)(V_xf + V_xi)t x_f = 500 + (1/2)(0 + 20)(5) x_f = 500 + 50 x_f = 550 therefore the total distance traveled is 550m Is this correct?

OpenStudy (australopithecus):

crud I forgot to calculate average velocity 550/35 = 16m/s thus V_xavg = 16m/s

OpenStudy (ujjwal):

Yep it seems ok. Total dist traveled = first part + second part + third part= 100+400+50=550 m Total time=10+20+5=35 sec.